# Dicke superposition probes for noise-resilient Heisenberg and super-Heisenberg Metrology

### v1

@@ II Parameter estimation and Fisher information
Parameter estimation is a core task underlying precision measurements in both classical and quantum metrology. Depending on the application, the parameter of interest may correspond to a frequency, phase, magnetic field, temperature, or other physical quantity to be estimated. Despite this diversity, any estimation protocol can be conceptually decomposed into three essential stages: (i) preparation of input probe, (ii) parameter encoding, and (iii) probe readout via optimal measurement and statistical inference.
@@ II.2 Quantum parameter estimation
where $\hat{\mathbf{J}}=(\hat{J}_{x},\hat{J}_{y},\hat{J}_{z})$ denotes the collective spin operator of an $N$-qubit system, with $\hat{J}_{\alpha}=\frac{1}{2}\,\sum_{i}\,\hat{\sigma}_{i\alpha}$, $\alpha=x,y,z$; $\hat{\sigma}_{i\alpha}$ denoting the Pauli matrices of the $i^{\rm th}$ qubit, and $\hat{\boldsymbol{n}}$ denotes a unit vector specifying the direction. The extremal eigenvalues of the Hamiltonian given by (12) are given by $\lambda_{\rm max}=N/2$ and $\lambda_{\rm min}=-N/2$, yielding
@@ II.3 Encoding with linear and nonlinear interactions
In contrast, genuinely entangled $N$-qubit probe states can achieve Heisenberg scaling [92, 66] of QFI under linear Hamiltonian encoding, with $F_{Q}(\hat{\rho}_{\rm ent},\hat{H})\propto N^{2}$ in which case enhanced precision of estimation is achieved:
@@ III Dicke superposition probes for noise-resilient linear metrology
Recent experiments have demonstrated remarkable control over multiqubit Dicke states across diverse physical platforms, including photonic systems, cold atoms, and trapped ions [56, 81, 82, 33, 87, 90, 79]. Furthermore, state-of-the-art experiments have achieved on-chip generation and coherent collective control of superpositions spanning the full Dicke manifold [17], employing nanofabricated silicon photonic circuits with programmable interferometers to prepare up to eight-photon Dicke superpositions $\left|D_{8-l,l}\right\rangle$ with $l=0,\ldots,8$.
@@ III.1 Linear metrology with Dicke superposition probes
To assess whether this limitation can be overcome, superpositions of two distinct Dicke states (excluding GHZ and W-superposition) are considered,
@@
For the $N$-qubit W–superposition state $\lvert\mathrm{W\bar{W}}\rangle~\equiv~\lvert D^{(N)}_{1,N-1}\rangle$, corresponding to an equal superposition of the single-excitation Dicke state and its spin-flipped counterpart, the QFI under encoding by the optimal Hamiltonian $\hat{H}_{1}=\hat{J}_{z}$ is $F_{Q}(\lvert D^{(N)}_{1,N-1}\rangle,\hat{J}_{z})=(N-2)^{2}$ for $N\geq 6$, which asymptotically approaches the Heisenberg-limit $F_{Q}(\lvert\mathrm{GHZ}\rangle,\hat{J}_{z})=N^{2}$ attained by the GHZ state [44].
@@ III.2 Noise resilience of near-optimal Dicke superposition states
Figures 2 and  3 compare the noise robustness of the near-optimal Dicke superposition state with that of the corresponding W–superposition, GHZ, and balanced Dicke states for $N=8$. Under phase damping, the Dicke superposition probe $\lvert D^{(8)}_{3,5}\rangle$ and the balanced Dicke state $\lvert D_{4,4}\rangle$ exhibit similar behavior, maintaining larger QFI over a broad range of noise strengths, outperform the GHZ and W–superposition probe $\lvert D^{(8)}_{1,7}\rangle$ (see Figs. 2(a), 3(a)). Under amplitude damping, the phase sensitivity of the balanced Dicke state lies between the SNL and the HL over a broader range of noise strengths than that of the GHZ, W–superposition, and near-optimal Dicke superposition states (see Fig. 3(b)). Under global depolarization, the optimal probe state (GHZ) and the near-optimal Dicke superposition state $|D^{(8)}_{3,5}\rangle$ exhibit similar metrological behavior, maintaining phase sensitivity between the HL and SNL across almost the entire range of noise strengths (see Fig. 3(c)).
@@ IV Metrology with collective two-body encoding Hamiltonians
Consider the following encoding Hamiltonians:
@@
Note that the eigenstates of the collective spin operator $\hat{J}_{\boldsymbol{n}}$ and its square $\hat{J}_{\boldsymbol{n}}^{\,2}$ are the Dicke states $\lvert D_{N-l,l}\rangle_{\boldsymbol{n}},l=0,1,\ldots,N$ with respect to the quantization axis $\hat{\boldsymbol{n}}$ and the associated eigenvalues given by (N/2)l(N/2)-l, $\left((N/2)-l\right)^{2}$ respectively. The maximum eigenvalue of $\hat{J}_{\boldsymbol{n}}$ is $N/2$, while that of $\hat{J}_{\boldsymbol{n}}^{\,2}$ is $N^{2}/4$. The latter is two-fold degenerate, with the corresponding eigenstates $\lvert D_{N,0}\rangle_{\boldsymbol{n}}=\lvert 0\rangle_{\boldsymbol{n}}^{\otimes N}$, and $\lvert D_{0,N}\rangle_{\boldsymbol{n}}=\lvert 1\rangle_{\boldsymbol{n}}^{\otimes N}.$ For the two-body encoding Hamiltonian $\hat{H}^{(1)}_{2}=\hat{J}_{\boldsymbol{n}}^{\,2}$, an equal-weight superposition of these two extremal eigenstates $\lvert\mathrm{GHZ}\rangle_{\boldsymbol{n}}=(\lvert 0\rangle_{\boldsymbol{n}}^{\otimes N}+\lvert 1\rangle_{\boldsymbol{n}}^{\otimes N})/\sqrt{2}$ is therefore an eigenstate associated with the maximum eigenvalue $N^{2}/4$.
@@
The eigenvalues $\lambda^{(r)}_{\rm max}$, $\lambda^{(r)}_{\rm min}$, the corresponding eigenvectors $|\phi^{(r)}_{\rm max}\rangle,|\phi^{(r)}_{\rm min}$ of $\hat{H}^{(r)}_{2}$, r=1,2,3,4r~=~1,2,3,4 along with the QFI $F_{Q}\!\left(\lvert\psi^{(r)}\rangle_{\rm opt},\,\hat{H}^{(r)}_{2}\right)$ are summarized in Table 2
@@ IV.1 Near-optimal Dicke superposition probes for nonlinear two-body encodings
Dicke superposition states $\lvert D^{(N)}_{l,l^{\prime}}\rangle$ were introduced in Sec. II in the context of linear (non-interacting) metrology, where they were shown to provide near-optimal phase sensitivity. It is natural to explore whether the same class of probes retains its metrological usefulness in the presence of nonlinear, two-body encoding Hamiltonians.
In the present context, the encoding dynamics is generated by collective two-body Hamiltonians $\hat{H}^{(r)}_{2}$, r=1,2,3,4r~=~1,2,3,4 (see Eqs.(IV)). The Fisher information is evaluated for Dicke superposition probes and optimized over the indices $(l,l^{\prime})$ for fixed system size. The resulting near-optimal Dicke superposition states and their corresponding Fisher information values for N=5N=588 are summarized in Table 4. Noise robustness of the optimal probe states and the near-optimal Dicke superposition states (encoded by the non-linear two-body interaction Hamiltonians $\hat{H}^{(r)}_{2}$r=1,2,3,4r~=~1,2,3,4) is investigated in the following subsection.
@@ IV.2 Noise resilience of optimal and near-optimal Dicke superposition probes in two-body metrology
Figure 5: Comparison of phase sensitivity $\Delta\theta$ of $N=8$-qubit optimal probe $\lvert\psi_{(N=8)}^{(2)}\rangle_{\boldsymbol{n}}=\dfrac{1}{\sqrt{2}}\left(\lvert{\rm GHZ}\rangle_{\boldsymbol{n}}+\lvert D_{4,4}\rangle_{\boldsymbol{n}}\right)$ with that of near optimal Dicke superposition probe $|D^{(8)}_{0,2}\rangle_{\boldsymbol{n}}$, both encoded via the two-body interaction Hamiltonian $\hat{H}^{(1)}_{2}=\hat{J}^{2}_{\boldsymbol{n}}$, subjected to global depolarization. The NL-SNL and the NL-HL are displayed for reference.
The impact of these three noise processes on the Fisher information of the identified probe states is summarized in Figures 4–7. Figure 4 displays the phase sensitivity of the optimal probe state $\lvert\psi_{(N=8)}^{(2)}\rangle_{\boldsymbol{n}}=\dfrac{1}{\sqrt{2}}\left(\lvert 0\rangle_{\boldsymbol{n}}^{\otimes 8}+\lvert D^{(8)}_{4,5}\rangle_{\boldsymbol{n}}\right)$ for $N=8$, subjected to three different noises after encoding the state using the Hamiltonian $\hat{H}^{(2)}_{2}=\hat{J}_{\boldsymbol{n}}+\hat{J}_{\boldsymbol{n}}^{2}$. It is observed, in general, that the optimal probes are particularly susceptible to local phase damping while exhibiting comparatively better resilience against amplitude damping and global depolarization, when they are encoded using two-body interaction Hamiltonians given in Eqs.(IV).
@@
Figures 5–7 present a direct comparison between the optimal probe states (see Eqs. (IV)) and the near-optimal Dicke superposition probes listed in Table 4. These results highlight noise regimes in which the Dicke superposition probes display metrological performance comparable to —and in certain cases outperforming — that of the corresponding optimal states under decoherence. In particular, Fig. 5 shows that both optimal and near-optimal Dicke superpositions respond similarly under global depolarizing noise, exhibiting comparable robustness for either class of probes in this case.
@@ V Conclusions
Building on the near-optimality of noise-resilient Dicke superposition probes in linear (non-interacting) metrology, a systematic extension to collective two-body interaction generators is carried out. For four representative two-body encoding Hamiltonians $\hat{H}^{(r)}_{2}$, $r=1,2,3,4$, the maximum and minimum eigenvalues, the corresponding eigenstates, and the resulting QFI are determined analytically, allowing the identification of optimal probe states in the noiseless setting. The associated scaling behavior confirms the emergence of nonlinear metrological limits, with separable probes bounded by a nonlinear SNL scaling and entangled probes capable of attaining the nonlinear HL scaling characteristic of two-body interactions. Near-optimal Dicke superposition probe states were identified by explicitly evaluating and optimizing the Fisher information.
@@
Note Added: During the final stages of preparing this manuscript, we became aware of a recent related work  [45] that investigates interaction-enhanced quantum metrology using specific maximal superpositions of symmetric Dicke states. While there is partial conceptual overlap in the use of Dicke superposition states, our work differs substantially in scope and focus. In particular, we identify near-optimal Dicke superposition probes and provide a unified analysis across both linear and two-body metrological schemes, with an explicit emphasis on noise resilience and robustness under realistic decoherence models.

### v2

@@ II Parameter estimation and Fisher information
Parameter estimation is a core task underlying precision measurements in both classical and quantum metrology. Depending on the application, the parameter of interest may correspond to a frequency, phase, magnetic field, temperature, or other physical quantity to be estimated. Despite this diversity, any estimation protocol can be conceptually decomposed into three essential stages: (i) preparation of input probe, (ii) parameter encoding, and (iii) probe readout via measurement and statistical inference.
@@ II.2 Quantum parameter estimation
where $\hat{\mathbf{J}}=(\hat{J}_{x},\hat{J}_{y},\hat{J}_{z})$ denotes the collective spin operator of an $N$-qubit system, with $\hat{J}_{\alpha}=\frac{1}{2}\,\sum_{i}\,\hat{\sigma}_{i\alpha}$, $\alpha=x,y,z$; $\hat{\sigma}_{i\alpha}$ denoting the Pauli matrices of the $i^{\rm th}$ qubit, and $\hat{\boldsymbol{n}}$ denotes a unit vector specifying the direction. The extremal eigenvalues of the Hamiltonian in (12) are given by $\lambda_{\rm max}=N/2$ and $\lambda_{\rm min}=-N/2$, yielding
@@ II.3 Encoding with linear and nonlinear interactions
In contrast, genuinely entangled $N$-qubit probe states can achieve Heisenberg scaling [92, 66] of QFI under linear Hamiltonian encoding, with $F_{Q}(\hat{\rho}_{\rm ent},\hat{H})\propto N^{2}$ in which case enhanced precision of estimation, the Heisenberg limit (HL) is achieved:
@@ III Dicke superposition probes for noise-resilient linear metrology
Recent experiments have demonstrated remarkable control over multiqubit Dicke states across diverse physical platforms, including photonic systems, cold atoms, and trapped ions [56, 81, 82, 33, 87, 90, 79]. Furthermore, state-of-the-art experiments have enabled on-chip generation and collectively coherent control of the superpositions of entire family of four-photon Dicke states [17] with the underlying architecture and control scheme extendable to more photons, providing a scalable route toward generating higher-photon-number Dicke superpositions.
@@ III.1 Linear metrology with Dicke superposition probes
To assess whether this limitation can be overcome, superpositions of two distinct Dicke states (excluding GHZ and W-superposition) are considered:
@@
For the $N$-qubit W–superposition state $\lvert\mathrm{W\bar{W}}\rangle~\equiv~\lvert D^{(N)}_{1,N-1}\rangle$, corresponding to an equal superposition of the single-excitation Dicke state and its spin-flipped counterpart, the QFI under encoding by the optimal Hamiltonian $\hat{H}_{1}=\hat{J}_{z}$ is $F_{Q}(\lvert D^{(N)}_{1,N-1}\rangle,\hat{J}_{z})=(N-2)^{2}$ for $N\geq 6$, and it asymptotically approaches the Heisenberg-limit $F_{Q}(\lvert\mathrm{GHZ}\rangle,\hat{J}_{z})=N^{2}$ attained by the GHZ state [44].
@@ III.2 Noise resilience of near-optimal Dicke superposition states
Figures 2 and  3 compare the noise robustness of the near-optimal Dicke superposition states with that of W–superposition, GHZ, and balanced Dicke states for $N=8$. Under phase damping, the Dicke superposition probe $\lvert D^{(8)}_{3,5}\rangle$ and the balanced Dicke state $\lvert D_{4,4}\rangle$ exhibit similar behavior, maintaining larger QFI over a broad range of noise strengths, outperform the GHZ and W–superposition probe $\lvert D^{(8)}_{1,7}\rangle$ (see Figs. 2(a), 3(a)). Under amplitude damping, the phase sensitivity of the balanced Dicke state lies between the SNL and the HL over a broader range of noise strengths than that of the GHZ, W–superposition, and near-optimal Dicke superposition states (see Fig. 3(b)). Under global depolarization, the optimal probe state (GHZ) and the near-optimal Dicke superposition state $|D^{(8)}_{3,5}\rangle$ exhibit similar metrological behavior, maintaining phase sensitivity between the HL and SNL across almost the entire range of noise strengths (see Fig. 3(c)).
@@ IV Metrology with collective two-body encoding Hamiltonians
Consider the following two-body interaction Hamiltonians:
@@
Note that the eigenstates of the collective spin operator $\hat{J}_{\boldsymbol{n}}$ and its square $\hat{J}_{\boldsymbol{n}}^{\,2}$ are the Dicke states $\lvert D_{N-l,l}\rangle_{\boldsymbol{n}},l=0,1,\ldots,N$ with respect to the quantization axis $\hat{\boldsymbol{n}}$ and the associated eigenvalues given by (N/2)l(N/2)-l, $\left((N/2)-l\right)^{2}$ respectively. The maximum eigenvalue of $\hat{J}_{\boldsymbol{n}}$ is $N/2$, while that of $\hat{J}_{\boldsymbol{n}}^{\,2}$ is $N^{2}/4$. The latter is two-fold degenerate, with the corresponding eigenstates $\lvert D_{N,0}\rangle_{\boldsymbol{n}}=\lvert 0\rangle_{\boldsymbol{n}}^{\otimes N}$, and $\lvert D_{0,N}\rangle_{\boldsymbol{n}}=\lvert 1\rangle_{\boldsymbol{n}}^{\otimes N}.$ For the two-body encoding Hamiltonian $\hat{H}^{(1)}_{2}=\hat{J}_{\boldsymbol{n}}^{\,2}$, an equal-weight superposition of these two extremal eigenstates is $\lvert\mathrm{GHZ}\rangle_{\boldsymbol{n}}=(\lvert 0\rangle_{\boldsymbol{n}}^{\otimes N}+\lvert 1\rangle_{\boldsymbol{n}}^{\otimes N})/\sqrt{2}$ is therefore an eigenstate associated with the maximum eigenvalue $N^{2}/4$.
@@
The eigenvalues $\lambda^{(r)}_{\rm max}$, $\lambda^{(r)}_{\rm min}$, the corresponding eigenvectors $|\phi^{(r)}_{\rm max}\rangle,|\phi^{(r)}_{\rm min}\rangle$ of $\hat{H}^{(r)}_{2}$, r=1,2,3,4r~=~1,2,3,4 along with the QFI $F_{Q}\!\left(\lvert\psi^{(r)}\rangle_{\rm opt},\,\hat{H}^{(r)}_{2}\right)$ are summarized in Table 2.
@@ IV.1 Near-optimal Dicke superposition probes for nonlinear two-body encodings
Having identified the specific class of Dicke superposition states that exhibit near-optimal phase sensitivity in linear (non-interacting) metrology, it is of natural interest to explore which class of states remains metrologically useful in two-body interaction metrology. In the present context, the encoding dynamics is generated by collective two-body Hamiltonians $\hat{H}^{(r)}_{2}$, r=1,2,3,4r~=~1,2,3,4 (see Eqs.(IV)). Upon optimizing over the direction $\boldsymbol{n}$, the Fisher information of Dicke superposition probes is evaluated, enabling the identification of near-optimal Dicke superposition states for a fixed system size. The resulting near-optimal states and their corresponding Fisher information values for N=5N=588 are summarized in Table 4. Noise robustness of the optimal probe states and the near-optimal Dicke superposition states (encoded by the non-linear two-body interaction Hamiltonians $\hat{H}^{(r)}_{2}$r=1,2,3,4r~=~1,2,3,4) is investigated in the following subsection.
@@ IV.2 Noise resilience of optimal and near-optimal Dicke superposition probes in two-body metrology
The impact of these three noise processes on the Fisher information of the identified probe states is summarized in Figures 4–7. Figure 4 displays the phase sensitivity of the optimal probe state $\lvert\psi_{(N=8)}^{(2)}\rangle_{\boldsymbol{n}}=\dfrac{1}{\sqrt{2}}\left(\lvert 0\rangle_{\boldsymbol{n}}^{\otimes 8}+\lvert D^{(8)}_{4,5}\rangle_{\boldsymbol{n}}\right)$ for $N=8$, subjected to three different noises, under encoding by the Hamiltonian $\hat{H}^{(2)}_{2}=\hat{J}_{\boldsymbol{n}}+\hat{J}_{\boldsymbol{n}}^{2}$. The nature of variation of phase sensitivity with the noise parameter $p$, exhibits similar nature for other three Hamiltonians $\hat{H}^{(r)}_{2}$, r=1, 3, 4r=1,\,3,\,4 also, It is observed, in general, that the optimal probes are particularly susceptible to local phase damping while exhibiting comparatively better resilience against amplitude damping and global depolarization, when they are encoded using two-body interaction Hamiltonians given in Eqs.(IV).
Figures 5–7 present a direct comparison between the optimal probe states (see Eqs. (IV)) and the near-optimal Dicke superposition probes listed in Table 4. These results highlight noise regimes in which the Dicke superposition probes display metrological performance comparable to —and in certain cases outperforming — that of the corresponding optimal states under decoherence. In particular, Fig. 5 shows that both optimal and near-optimal Dicke superpositions respond similarly under global depolarizing noise, exhibiting comparable robustness in both classes of probes, whereas Fig. 6 illustrates that near-optimal probes outperform optimal ones under dephasing noise.
@@
Figure 5: Comparison of phase sensitivity $\Delta\theta$ of $N=8$-qubit optimal probe $\lvert\psi_{(N=8)}^{(2)}\rangle_{\boldsymbol{n}}=\dfrac{1}{\sqrt{2}}\left(\lvert{\rm GHZ}\rangle_{\boldsymbol{n}}+\lvert D_{4,4}\rangle_{\boldsymbol{n}}\right)$ with that of near optimal Dicke superposition probe $|D^{(8)}_{0,2}\rangle_{\boldsymbol{n}}$, both encoded via the two-body interaction Hamiltonian $\hat{H}^{(1)}_{2}=\hat{J}^{2}_{\boldsymbol{n}}$, subjected to global depolarization. The NL-SNL and the NL-HL are displayed for reference. The optimal and near-optimal states corresponding to other three two-body interaction Hamiltonians $\hat{H}^{(r)}_{2}$, $r=2,3,4$ exhibit analogous response to global depolarizing noise.
@@ V Conclusions
Building on the near-optimality of noise-resilient Dicke superposition probes in linear (non-interacting) metrology, a systematic extension to collective two-body interaction generators is carried out. For four representative two-body encoding Hamiltonians $\hat{H}^{(r)}_{2}$, $r=1,2,3,4$, the maximum and minimum eigenvalues, the corresponding eigenstates, and the resulting QFI are determined analytically, allowing for the identification of optimal probe states in the noiseless setting. The associated scaling behavior confirms the emergence of nonlinear metrological limits, with separable probes bounded by a nonlinear SNL scaling and entangled probes capable of attaining the nonlinear HL scaling characteristic of two-body interactions. Near-optimal Dicke superposition probe states were identified by explicitly evaluating and optimizing the Fisher information.
@@
Note Added: During the final stages of preparing this manuscript, we became aware of a recent related work  [45] that investigates interaction-enhanced quantum metrology using specific maximal superpositions of symmetric Dicke states. While there is partial overlap in the near-optimal Dicke superposition states of linear metrology, our work differs substantially in scope and focus. In particular, we provide a unified analysis across both linear and two-body metrological schemes, with an explicit emphasis on noise resilience and robustness of optimal and near-optimal states under realistic decoherence models.